OpenStudy (anonymous):

A building measures 50 feet in length and w feet in width. Which of the following is an expression for the perimeter of the building? A. 50/w B. 50w C. 50 + w D. 100 + 2w

OpenStudy (anonymous):

D

OpenStudy (anonymous):

 Perimeter formula of rectangle is.. P=a+b+a+b P=2a+2b in your case you have.. P=2*50+2w P=100+2w